amayawarren:

If the sphere shown above has a radius of 19 units, then what is the approximate volume of the sphere? A. 9,145.33 cubic units B. 1,444 cubic units C. 11,431.67 cubic units D. 4,572.67 cubic units

AZ:

Volume of a sphere = \(\dfrac{4}{3}\pi r^3\) where r is your radius and they told you r = 19 so plug that in and solve for the volume

AZ:

volume = \(\dfrac{4}{3} \times \pi \times (19)^3\) volume = ??